OpenStudy (anonymous):

Algebra....Soving Equations with Radicals The square root of 3x +10 = 2x-5 ???

OpenStudy (helder_edwin):

\[ \large \sqrt{3x+10}=2x-5 \] ??

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Edwin???

OpenStudy (helder_edwin):

square each side of the equation

OpenStudy (anonymous):

\[3x+10=4x ^{2}-20x+25\]

OpenStudy (helder_edwin):

yes now it is a quadratic equation can u solve this?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

I'm stuck at this point

OpenStudy (helder_edwin):

put everything on one side \[ \large 0=4x^2-20x+25-3x-10=4x^2-23x+15 \]

OpenStudy (anonymous):

where did the 25 come from?

OpenStudy (helder_edwin):

now \[ \large 0=4x^2-23x+15=\frac{(4x-20)(4x-3)}{4} \] \[ \large =\frac{4(x-5)(4x-3)}{4}=(x-5)(4x-3) \]

OpenStudy (helder_edwin):

u posted it

OpenStudy (anonymous):

oh. O.K.

OpenStudy (anonymous):

x=5, x=3/4 ???

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

did u understand it?

OpenStudy (anonymous):

yes....Thanks Edwin....!!!

OpenStudy (helder_edwin):

u r welcome my name is Helder (Edwin is my middle name)